# Bayesian Optimisation with Unknown Hyperparameters: Regret Bounds Logarithmically Closer to Optimal

### v1

@@ 7 Conclusions
Within this work, we addressed the problem of BO with unknown hyperparameters. We proposed an algorithm with a cumulative regret bound logarithmically closer to optimal than the previous state of the art and showed that our algorithm can outperform existing baselines in practice. One limitation of our work is that we only showed how to handle the isotropic case, i.e. where the same length scale value is applied for every dimension. This limitation is because our algorithm requires the knowledge of the regret bounds of an optimiser utilising a given length scale value, which in turn requires the knowledge of the bounds on MIG $\gamma(k^{\theta})$ for the used kernel. To the best of our knowledge, within the existing literature, no work has yet derived those bounds in non-isotropic cases. However, we believe that if such bounds were obtained, one could easily extend our algorithm to the non-isotropic case, in the same way as we extended our base algorithm to handle unknown norm and length scale simultaneously. This constitutes a promising direction of future work.
Appendix A General Hyperparameter case
@@
11:        $U_{t+1}=\left\{u\in U_{t}:L_{t}(u)+\frac{1}{|S_{t}^{u}|}\sum_{\tau\in S_{t}^{u}}\beta_{\tau}^{u}\sigma^{u}_{\tau-1}(\bm{x}_{\tau})\geq\max_{u^{\prime}\in U_{t}}L_{t}(u^{\prime})\right\}$
@@
We first derive a general result for a BO Algorithm utilising a regret-balancing scheme, under hyperparameters of any form. We assume the unknown hyperparameter takes values in some set $\mathcal{U}$ and specifying this hyperparameter uniquely defines a GP model that conditioned on data provides us with mean and variance we can use to construct UCB acquisition function.
Lemma A.1.
Let us run Algorithm 2 with a given choice of the hyperparameter-proposing function $a:\mathbb{N}\to\mathcal{U}$ and denote by $A=\bigcup_{t=0,\dots,T}a(t)$ to be the set of all hyperparameters introduced by $a(\cdot)$. Let $t_{0}$ be the first iteration, where a well-specified candidate is introduced and $\mathcal{T}$ be the set of all iterations after $t_{0}$, where at least one hyperparameter is rejected by operation in line 11. We then have that:
$\sum_{t\notin\mathcal{T}}r_{t}=\mathcal{O}\left(|A|B+\left(R^{u^{*}}(T)+\sqrt{T\xi_{T}}\right)\left(\sqrt{\sum_{u\in A}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}+|A|\right)\right),$
where $u^{\star}=\arg\min_{u\in a(t_{0})}R^{u}(T)$ and $\xi_{t}=2\sigma_{N}^{2}\log\frac{|A|\pi^{2}t^{2}}{6\delta_{B}}$.
Proof.
We first recall a result from the literature.
Lemma A.2 (Lemma 5.1 in [36]).
@@
We will prove the bound assuming the events of Theorem 2.1 and Lemma A.2 hold. As this happens with probability at least $1-\delta_{A}-\delta_{B}$, the bound also holds with the same probability.
Optimal hyperparameter $u^{*}$ is never removed First, observe that under these events, the true hyperparameter $u^{*}$ will never be removed. To see this, observe that:
$\frac{1}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)=\frac{1}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\left(f(\bm{x}^{*})-y_{t}-\epsilon_{t})\right)\leq\frac{R_{u}(|S_{t}^{u^{*}}|)}{S_{t}^{u^{*}}}$
@@
$f(\bm{x}^{*})$ $\leq\frac{R_{u}(|S_{t}^{u^{*}}|)}{S_{t}^{u^{*}}}+\frac{1}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\epsilon_{t}+\sum_{t\in S_{t}^{u^{*}}}\frac{y_{t}}{|S_{t}^{u^{*}}|}$
$\leq\frac{R_{u}(|S_{t}^{u^{*}}|)}{S_{t}^{u^{*}}}+\sqrt{\frac{\xi_{t}}{|S_{t}^{u}|}}+\sum_{t\in S_{t}^{u^{*}}}\frac{y_{t}}{|S_{t}^{u^{*}}|}.$ (1)
Now, for any $u\in A$ and any $t\in S_{T}^{u}$, we have the following:
@@
$=\sum_{t\in S_{t}^{u}}\frac{y_{t}}{|S_{t}^{u}|}-\sum_{t\in S_{t}^{u}}\frac{\epsilon_{t}}{|S_{t}^{u}|}\leq\sum_{t\in S_{t}^{u}}\frac{y_{t}}{|S_{t}^{u}|}-\sqrt{\frac{\xi_{t}}{|S_{t}^{u}|}}.$ (2)
Thus combining Equations A and A, we get that for any $u\in U$:
$\frac{R_{u}(|S_{t}^{u^{*}}|)}{S_{t}^{u^{*}}}+\sqrt{\frac{\xi_{t}}{|S_{t}^{u}|}}+\sum_{t\in S_{t}^{u^{*}}}\frac{y_{t}}{|S_{t}^{u^{*}}|}\geq\sum_{t\in S_{t}^{u}}\frac{y_{t}}{|S_{t}^{u}|}-\sqrt{\frac{\xi_{t}}{|S_{t}^{u}|}},$
which means the if statement in line 11 will always evaluate to true and thus $u^{*}\in U_{t}$ for all $t=1,\dots,T$. ∎
Bound on $\sum_{t\notin\mathcal{T}}r_{t}$ Now, notice that on all iterations $t\notin\mathcal{T}$, all hyperparameters in $U_{t}$ must satisfy the if statement in line 11. Thus:
$\frac{1}{|S_{t}^{u}|}\sum_{t\in S_{t}^{u}}y_{t}+\frac{R_{u}(|S_{t}^{u}|)}{|S_{t}^{u}|}+\sqrt{\frac{\xi_{t}}{|S_{t}^{u}|}}\geq\frac{1}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}y_{t}-\sqrt{\frac{\xi_{t}}{|S_{t}^{u^{*}}|}}.$
@@
$\frac{1}{|S_{t}^{u}|}\sum_{t\in S_{t}^{u}}\left(f(\bm{x}^{*})-y_{t}\right)-\frac{R_{u}(|S_{t}^{u}|)}{|S_{t}^{u}|}-\sqrt{\frac{\xi_{t}}{|S_{t}^{u}|}}\leq\frac{1}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\left(f(\bm{x}^{*})-y_{t}\right)+\sqrt{\frac{\xi_{t}}{|S_{t}^{u^{*}}|}}.$
Using the fact that $y_{t}=f(\bm{x}_{t})+\epsilon_{t}$ and Lemma A.2 we get:
$\frac{1}{|S_{t}^{u}|}\sum_{t\in S_{t}^{u}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)-\frac{R_{u}(|S_{t}^{u}|)}{|S_{t}^{u}|}-2\sqrt{\frac{\xi_{t}}{|S_{t}^{u}|}}\leq\frac{1}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)+2\sqrt{\frac{\xi_{t}}{|S_{t}^{u^{*}}|}}.$
Rearranging, we obtain:
$\sum_{t\in S_{t}^{u}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)\leq R_{u}(|S_{t}^{u}|)+2\sqrt{|S_{t}^{u}|\xi_{t}}+\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)+2|S_{t}^{u}|\sqrt{\frac{\xi_{t}}{|S_{t}^{u^{*}}|}}.$
We now recall the balancing condition, which is satisfied by regret balancing algorithms as proven in Lemma 5.2 of [29]. In their proof, the maximum possible regret value was 1, we thus need to rescale it to our setting, where this maximum value is $2B$. Applying this fact to the previous display we get:
$\sum_{t\in S_{t}^{u}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)\leq R^{u^{*}}(|S_{t}^{u^{*}}|)+2B+2\sqrt{|S_{t}^{u}|\xi_{t}}+\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)+2|S_{t}^{u}|\sqrt{\frac{\xi_{t}}{|S_{t}^{u^{*}}|}}.$
Plugging this into the cumulative regret bound, we get:
@@
$\sum_{t\notin\mathcal{T}}r_{t}$ $=\sum_{u\in U}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{t}^{u}\end{subarray}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)$
$\leq\sum_{u\in U}\left(R^{u^{*}}(|S_{t}^{u^{*}}|)\left(1+\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}\right)+2B+2\sqrt{|S_{t}^{u}|\xi_{t}}+2\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\sqrt{|S_{t}^{u}|\xi_{t}}\right)$
$=R^{u^{*}}(|S_{t}^{u^{*}}|)\left(|A|+\sum_{u\in U}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}\right)+|A|B+2\sum_{u\in U}\sqrt{|S_{t}^{u}|\xi_{t}}+\sum_{u\in U}\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\sqrt{|S_{t}^{u}|\xi_{t}}$
$\leq R^{u^{*}}(|S_{t}^{u^{*}}|)\left(|A|+\sum_{u\in U}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}\right)+|A|B+2\sqrt{|A|T\xi_{T}}+\sqrt{\sum_{u\in U}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\sqrt{T\xi_{t}}$
$=\mathcal{O}\left(R^{u^{*}}(|S_{t}^{u^{*}}|)\left(|A|+\sum_{u\in U}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}\right)+|A|B+\left(\sqrt{|A|}+\sqrt{\sum_{u\in U}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\right)\sqrt{T\xi_{t}}\right),$
where the penultimate transition is due to Cauchy-Schwarz. Now, since the suspected regret bounds in BO will be of form $R^{u}(T)=\sqrt{T\beta^{u}_{T}\gamma_{T}^{u}}(\sqrt{\gamma_{T}^{u}}+\sqrt{B})$, we have that $R^{u}(T)=\mathcal{O}(\gamma_{T}^{u}\sqrt{T\beta^{u}_{T}})$. Substituting this fact, we get:
$R^{u^{*}}(|S_{t}^{u^{*}}|)\sum_{u\in U}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}$ $\leq\mathcal{O}\left(\sum_{u\in U}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}\gamma_{T}^{u}\sqrt{|S_{t}^{u^{*}}|\beta^{u}_{T}}\right)$
$\leq\mathcal{O}\left(\sum_{u\in U}\sqrt{\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\gamma_{T}^{u}\sqrt{|S_{t}^{u}|\beta^{u}_{T}}\right)$
$\leq\mathcal{O}\left(\sqrt{\sum_{u\in U}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}\gamma_{T}^{u}\sqrt{T\beta^{u}_{T}}\right)$
$\leq\mathcal{O}\left(\sqrt{\sum_{u\in U}\frac{|S_{t}^{u}|}{|S_{t}^{u^{*}}|}}R^{u}(T)\right)$
We thus get the final bound of the form:

### v3

@@ 7 Conclusions
Within this work, we addressed the problem of BO with unknown hyperparameters. We proposed an algorithm with a cumulative regret bound logarithmically closer to optimal than the previous state of the art and showed that our algorithm can outperform existing baselines in practice. One limitation of our work is that we only showed how to handle the isotropic case, i.e. where the same length scale value is applied for every dimension. This limitation is because our algorithm requires the knowledge of the regret bounds of an optimiser utilising a given length scale value, which in turn requires the knowledge of the bounds on MIG $\gamma_{T}(k^{\theta})$ for the used kernel. To the best of our knowledge, within the existing literature, no work has yet derived those bounds in non-isotropic cases. However, we believe that if such bounds were obtained, one could easily extend our algorithm to the non-isotropic case, in the same way as we extended our base algorithm to handle unknown norm and length scale simultaneously. This constitutes a promising direction of future work.
Another limitation of our work is the assumption of known noise magnitude $\sigma_{N}$. The problem of simultaneously not knowing kernel hyperparameters and noise magnitude is extremely challenging, as large variations in observed function values can be a result of either short lengthscale value or large noise magnitude. To the best of our knowledge, previous work did not tackle this setting and it remains an open problem.
Our algorithm relies on the standard GP model, which can result in poor scalability to large datasets and high-dimensional spaces. Extending our work to sparse GPs [29, 31] and kernels specifically designed for a high number of dimensions [14, 42] is another possible direction of future work.
Acknowledgments and Disclosure of Funding
We would like to thank Toni Karvonen for pointing out important properties of the $\nu$-Matérn kernel’s RKHS, which were crucial to take into account in the proposed problem setting. We would also like to thank Ondrej Bajgar and the anonymous reviewers for their helpful comments about improving the paper. Juliusz Ziomek was supported by the Oxford Ashton-Memorial Scholarship and EPSRC DTP grant EP/W524311/1. Masaki Adachi was supported by the Clarendon Fund, the Oxford Kobe Scholarship, the Watanabe Foundation, and Toyota Motor Corporation.
Appendix A Proof of Proposition 2.2
See 2.2
Proof.
The RBF case follows directly from Proposition 2 in [6]. The same Proposition also provides a bound for the $\nu$-Matérn case, but more recent results allow us to derive a tighter bound. Section B.2. in [6] proves that the $\nu$-Matérn kernel has the ($C_{p}$,$\beta_{P}$) polynomial eigendecay: $\beta_{P}=(2\nu+d)/d$ and $C_{p}=\left(\frac{1}{\theta}\right)^{2\nu+d}$, according to the definition of polynomial eigendecay as given by [37]. Substituting these values to Corollary 1 of [37], we get that in $\nu$-Matérn case
$\mathcal{I}_{T}^{\theta}=\mathcal{O}\left(C_{p}^{1/\beta_{P}}T^{1/\beta_{P}}\log^{1-1/\beta_{P}}(T)^{1-1/\beta_{P}}\right)=\mathcal{O}\left(\left(\frac{1}{\theta}\right)^{d}T^{d/(2\nu+d)}\log^{2\nu/(d+2\nu)}(T)^{2\nu/(d+2\nu)}\right),$
which finishes the proof. ∎
Appendix B General Hyperparameter case
Lemma B.1Theorem 2.1Lemma B.2Lemma B.3Lemma B.4Lemma B.5
Figure 5: Diagram of the relationship between Lemmas in this Section. An incoming arrow means that the Lemma relies on the Lemma/ Theorem from which the arrow is outgoing. The final objective of this Section is to prove Lemma B.5.
We first derive a general result for a BO Algorithm utilising a regret-balancing scheme, under hyperparameters of any form. We present the pseudo-code of that procedure in Algorithm 2. We now proceed the prove the general regret bound for this algorithm. We would like to note that our proof closely follows the idea of [33]. We assume the unknown hyperparameter takes values in some set $\mathcal{U}$ and specifying this hyperparameter uniquely defines a kernel function $k^{u}(\cdot,\cdot)$ and the RKHS norm $B^{u}$. We say a hyperparameter value is well-specified if $|f|_{k^{u}}\leq B^{u}$. Algorithm 2 requires hyperparameter-proposing function $a(\cdot)$ as one of the inputs, which optionally expands the set of considered hyperparameters. We will denote by $A=\bigcup_{t=0,\dots,T}a(t)$ the set of all hyperparameters introduced by $a(\cdot)$ up to step $T$. We will also denote $\mathcal{W}$ to mean the set of all well-specified hyperparameters in $A$ and $\mathcal{M}=A\setminus\mathcal{W}$ to denote the set of misspecified hyperparameters. Let $u^{\star}$ be the first well-specified hyperparameter introduced. We will also write $\mathcal{T}$ to mean the set of all iterations after (and including) $t_{0}$, where at least one hyperparameter value was rejected, that is:
$\mathcal{T}=\left\{t=t_{0},\dots,T\Bigg{|}\exists_{u\in U_{t}}L_{t}(u)+\frac{2}{|S_{t}^{u}|}\sum_{\tau\in S_{t}^{u}}\beta_{\tau}^{u}\sigma^{u}_{\tau-1}(\bm{x}_{\tau})<\max_{u^{\prime}\in U_{t}}L_{t}(u^{\prime})\right\}.$
We also define $\mathcal{M}_{0}=\{u\in\mathcal{M}|\exists_{t\geq t_{0}}u_{t}=u\}$ to be the set of all misspecified hyperparameters that were selected at least once after $t_{0}$.
@@ Appendix B General Hyperparameter case
11:        $U_{t+1}=\left\{u\in U_{t}:L_{t}(u)+\frac{2}{|S_{t}^{u}|}\sum_{\tau\in S_{t}^{u}}\beta_{\tau}^{u}\sigma^{u}_{\tau-1}(\bm{x}_{\tau})\geq\max_{u^{\prime}\in U_{t}}L_{t}(u^{\prime})\right\}$
@@
We will first provide some auxiliary Lemmas that we will be required to prove the general result for Algorithm 2. We first recall a result from existing literature on the concentration of noise.
Lemma B.1 (Lemma 5.1 in [41]).
@@
The next result we will need is the high-probability guarantee that the optimal hyperparameter $u^{\star}$ will not be removed from the set of considered hyperparameters after it is introduced.
Lemma B.2.
If the events of Theorem 2.1 and Lemma B.1 hold, then the optimal hyperparameter $u^{\star}$ is never removed after it is introduced, i.e. $u^{*}\in U_{t}$ for all $t=t_{0},\dots,T$, where $t_{0}$ is the first iteration, where the optimal hyperparameter $u^{\star}$ is introduced.
Proof.
First, observe that under these events, the optimal hyperparameter $u^{*}$ will never be removed. To see this, observe that at any time set $t$ such that $u_{t}=u^{\star}$ the following holds:
$f(\bm{x}^{*})-f(\bm{x}_{t})$ $\leq\mu_{t}^{u^{\star}}(\bm{x}^{\star})+\beta_{t}^{u^{\star}}\sigma_{t}^{u^{\star}}(\bm{x}^{\star})-\mu_{t}^{u^{\star}}(\bm{x}_{t})+\beta_{t}^{u^{\star}}\sigma_{t}^{u^{\star}}(\bm{x}_{t})$
$\leq\mu_{t}^{u^{\star}}(\bm{x}_{t})+\beta_{t}^{u^{\star}}\sigma_{t}^{u^{\star}}(\bm{x}_{t})-\mu_{t}^{u^{\star}}(\bm{x}_{t})+\beta_{t}^{u^{\star}}\sigma_{t}^{u^{\star}}(\bm{x}_{t})$
$=2\beta_{t}^{u^{\star}}\sigma_{t}^{u^{\star}}(\bm{x}_{t}),$
where the first inequality follows from the fact that the event of Theorem 2.1 holds and $u^{\star}$ is well-specified, and the second inequality is because we chose $\bm{x}_{t}$ so as to maximise the UCB. Using this fact, we get that:
$\frac{1}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\left(f(\bm{x}^{*})-(y_{t}-\epsilon_{t})\right)=\frac{1}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\left(f(\bm{x}^{*})-f(\bm{x}_{t})\right)\leq\frac{2}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\beta_{t}^{u^{\star}}\sigma_{t}^{u^{\star}}(\bm{x}_{t})$
@@
$f(\bm{x}^{*})$ $\leq\frac{2}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\beta_{t}^{u^{\star}}\sigma_{t}^{u^{\star}}(\bm{x}_{t})-\frac{1}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\epsilon_{t}+\sum_{t\in S_{t}^{u^{*}}}\frac{y_{t}}{|S_{t}^{u^{*}}|}$
$\leq\frac{2}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\beta_{t}^{u^{\star}}\sigma_{t}^{u^{\star}}(\bm{x}_{t})+\sqrt{\frac{\xi_{t}}{|S_{t}^{u^{\star}}|}}+\sum_{t\in S_{t}^{u^{*}}}\frac{y_{t}}{|S_{t}^{u^{*}}|},$ (1)
where the second inequality is a result of Lemma B.1. Now, for any $u\in A$ and any $t\in S_{T}^{u}$, we have the following:
@@
$=\sum_{t\in S_{t}^{u}}\frac{y_{t}}{|S_{t}^{u}|}-\sum_{t\in S_{t}^{u}}\frac{\epsilon_{t}}{|S_{t}^{u}|}\geq\sum_{t\in S_{t}^{u}}\frac{y_{t}}{|S_{t}^{u}|}-\sqrt{\frac{\xi_{t}}{|S_{t}^{u}|}},$ (2)
where again second inequality comes from Lemma B.1. Thus combining Inequalities B and B, we get that for any $u\in A$:
$\frac{2}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}\beta_{t}^{u^{\star}}\sigma_{t}^{u^{\star}}(\bm{x}_{t})+\sqrt{\frac{\xi_{t}}{|S_{t}^{u^{\star}}|}}+\sum_{t\in S_{t}^{u^{*}}}\frac{y_{t}}{|S_{t}^{u^{*}}|}\geq\sum_{t\in S_{t}^{u}}\frac{y_{t}}{|S_{t}^{u}|}-\sqrt{\frac{\xi_{t}}{|S_{t}^{u}|}},$
which means the if statement in line 11 will always evaluate to true for $u^{\star}$ and thus $u^{*}\in U_{t}$ for all $t=1,\dots,T$. ∎
We now recall the balancing condition, which is satisfied by regret balancing algorithms as proven in Lemma 5.2 of [33]. This condition basically says that due to the selection rule that chooses an algorithm with current lower suspected regret, the regret of any two algorithms (that have not been rejected) must be "close" to each other. We prove a slightly different statement than that in [33], as in our case the set of base algorithms can be dynamically expanded.
Lemma B.3.
Assume the event of Lemma B.2 holds. For any $u\in A$ that was selected in line 3 at least once after $t_{0}$ and any time step $t=t_{0},\dots,T$ we must have that $R^{u}(|S_{t}^{u}|)\leq R^{u^{\star}}(|S_{t}^{u^{\star}}|)+2B^{u^{\star}}$.
Proof.
We will prove the statement by contradiction. If the statement of the Lemma was not true, we would have:
$R^{u}(|S_{t}^{u}|)>R^{u^{\star}}(|S_{t}^{u^{\star}}|)+2B^{u^{\star}}\geq R^{u^{\star}}(|S_{t}^{u^{\star}}|+1),$
where the second inequality comes from the assumption that the bounds are non-trivial. If $u$ has been selected at least once after $u^{\star}$ was introduced (and by Lemma B.2 $u^{\star}$ could not have been excluded afterwards), then the last time $u$ was selected, $u^{\star}$ must have been present in $U_{t}$ and as such selection of $u$ such that $R^{u}(|S_{t}^{u}|)>R^{u^{\star}}(|S_{t}^{u^{\star}}|+1)$ would violate the selection rule in line 3. ∎
Before we can prove the final result of this Section, we need one more auxiliary Lemma. This Lemma bounds the regret on all the iterations, where none of the hyperparameters was rejected (i.e. $t\notin\mathcal{T}$), by using the rejection rule of line 11.
Lemma B.4.
Assume the event of Lemma B.3 holds. Let us call by $\mathcal{T}$ the set of all iterations after $t_{0}$, where we do not reject any of the candidates. We must have that for any $u\in A$:
$\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{T}^{u}\end{subarray}}r_{t}$ $\leq\left(\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}+1\right)CR^{u^{\star}}(|S_{T}^{u^{*}}|)+2CB^{u^{\star}}+2\left(\sqrt{\frac{|S_{T}^{u}|}{|S_{T}^{u^{*}}|}}+1\right)\sqrt{|S_{T}^{u}|\xi_{t}}.$
Proof.
Let $t^{\prime}$ be the smallest iteration after $t_{0}$, such that $S_{t^{\prime}}^{u}=S_{T}^{u}$ and $t^{\prime}\notin\mathcal{T}$, i.e. $t^{\prime}$ is the last iteration not in $\mathcal{T}$ when $u$ was played. Notice that since $t^{\prime}\notin\mathcal{T}$, all hyperparameter values in $U_{t^{\prime}}$ must satisfy the if statement in line 11 when compared with any other hyperparameter value in $U_{t^{\prime}}$. Let us choose any hyperparameter value in $U_{t^{\prime}}$ and compare it with $u^{\star}$ (which must be in $U_{t^{\prime}}$ due to the event of Lemma B.3 that guarantees preservation of $u^{\star}$.). For the if statement to evaluate to true, we must have:
$\frac{1}{|S_{t^{\prime}}^{u}|}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{t^{\prime}}^{u}\end{subarray}}y_{t}+\frac{2}{|S_{t^{\prime}}^{u}|}\sum_{t\in S_{t^{\prime}}^{u}}\beta_{t}^{u}\sigma_{t}^{u}(\bm{x}_{t})+\sqrt{\frac{\xi_{t}}{|S_{t^{\prime}}^{u}|}}\geq\frac{1}{|S_{t}^{u^{*}}|}\sum_{t\in S_{t}^{u^{*}}}y_{t}-\sqrt{\frac{\xi_{t}}{|S_{t}^{u^{*}}|}}.$
@@
$\frac{1}{|S_{t^{\prime}}^{u}|}\sum_{\begin{subarray}{c}t\notin\mathcal{T}\\ t\in S_{t^{\prime}}^{u}\end{subarray}}\left(f(\bm{x}^{*})-y_{t}\right)-\frac{2\sum_{t\in S_{t^{\prime}}^{u}}\beta_{t}^{u}\sigma_{t}^{u}(\bm{x}_{t})}{|S_{t^{\prime}}^{u}|}-\sqrt{\frac{\xi_{t}}{|S_{t^{\prime}}^{u}|}}\leq\frac{1}{|S_{t^{\prime}}^{u^{\star}}|}\sum_{t\in S_{t^{\prime}}^{u^{\star}}}\left(f(\bm{x}^{*})-y_{t}\right)+\sqrt{\frac{\xi_{t}}{|S_{t^{\prime}}^{u^{\star}}|}}.$
Using the fact that $y_{t}=f(\bm{x}_{t})+\epsilon_{t}$ and Lemma B.1 we get that: